The tension at which a fishing line snaps is commonly called the line's “strength.” What minimum strength is needed for a line that is to stop a salmon of weight 93 N in 12 cm if the fish is initially drifting horizontally at 2.7 m/s? Assume a constant deceleration.
Answer #1

Step 1:

First we need to find deceleration of fish to stop it completely

Using 3rd kinematic equation:

V^2 = U^2 + 2*a*d

U = Initial speed of fish = 2.7 m/sec

V = final speed of fish = 0 m/sec

d = distance at which fish needs to be stopped = 12 cm = 0.12 m

So,

a = (V^2 - U^2)/(2*d)

a = (0^2 - 2.7^2)/(2*0.12)

a = deceleration = -30.375 m/sec^2

Now given that weight of Salmon = 93 N

m = Mass of Salmon = Weight/g = 93/9.81 = 9.48 kg

So Now minimum force applied on Salmon due to fish will be:

Using Newton's 2nd law:

T = m*a

T = 9.48*30.375 = 287.955 N

T = 288 N = Minimum Strength of wire

Let me know if you've any query.